Let f be a function defined on R by:


F(x)= { x+5/(x^2-25) , when x≠5


{ 1 when x= 5


Check whether f is uniformly continuous on [-3,3] or not

1
Expert's answer
2022-03-01T12:08:09-0500

ANSWER. The function F is uniformly continuous on [-3,3]

EXPLANATION.

"F(x) =\left\{\begin{matrix}
\frac{x+5}{ x^{2}-25} ,&if\, x\neq 5 \\ 
 1,& if x=5 
\end{matrix}\right.=\left\{\begin{matrix}
\frac{x+5}{ (x-5)(x+5) } ,&if\, x\neq 5 \\ 
 1,& if x=5 
\end{matrix}\right." . The function "F" is defined on the set "D=\left ( -\infty,-5 \right )\cup\left ( -5,+\infty \right )" . If "x\in[-3,3]" , then "F(x)=\frac{1 }{x-5}" and {5} "\notin [-3,3]" .



Therefore, on the segment [-3,3] the function "F" is rational and is defined at all points of the segment. So, "F" is continuous (on [-3,3]). By Cantor's Theorem , a function continuous on a bounded and closed set [-3,3] is a uniformly continuous on this set.